# GreenHyperSpectra: A multi-source hyperspectral dataset for global vegetation trait prediction

### v1

@@ 3 The GreenHyperSpectra dataset
We curated spectral data from multiple instruments across three primary platforms: proximal, airborne, and spaceborne (Figure 2). All data were processed to the level of at-surface reflectance. Proximal measurements were obtained using field spectrometers such as the ASD FieldSpec and SVC HR-1024i, typically positioned in a close range in nadir orientation to record top-of-canopy reflectance. Airborne data were acquired using high-resolution sensors, including the AVIRIS-Next Generation, AVIRIS-Classic, NEON Airborne Observation Platform (AOP) and Specim AISAFenix instruments, which cover landscape-level vegetation scenes with variable viewing geometries and meter-scale spatial resolutions. Space-borne acquisitions were collected from missions such as PRISMA, Hyperion, EMIT, and EnMAP, offering a larger scale
@@
The dataset includes samples from diverse biomes, with acquisitions spanning from 1992 to 2024, capturing broad ecological, climatic, and seasonal variability across a wide range of environments. Figure 2 maps the global spatial distribution of GreenHyperSpectra and a pool of previously aggregated datasets for plant trait prediction (see details in § 4 ). While the compiled labeled dataset is spatially limited, GreenHyperSpectra encompasses substantially broader spatial coverage and environmental heterogeneity, better representing real-world remote sensing operational conditions (more details in Appendix A).
@@ 5 Experimental settings
We conduct comprehensive ablation experiments across several dimensions to consistently evaluate design trade-offs of the MAE models in spectral representation learning. First, we explore architectural complexity through a grid search spanning transformer configurations with varying numbers of layers $\{6,8,10\}$ and attention heads $\{4,8,16\}$. We select the configuration demonstrating optimal performance on the downstream trait prediction task for subsequent experiments. Considering this optimal architecture, we investigate alternative loss formulations for spectral reconstruction. Beyond conventional MSE, we examine hybrid approaches which incorporate cosine similarity loss weighted by a coefficient $\alpha\in\{1,0.1,0.01,0\}$ to enhance capture of spectral shape specificity. Finally, we assess the effect of token granularity by varying patch sizes (10, 20, 40, and 430) used during spectral masking and reconstruction in the full-range scenario. These targeted ablations informed the final MAE configuration used in our other benchmark evaluations. All results of the aforementioned ablation studies are presented in Tables 14, 15 and 16 in Appendix B.3.
@@
For each experimental setting, we report the performance metrics averaged across three random seeds to measure the variability related to stochastic training effects. Our evaluation framework employs two complementary metrics: the coefficient of determination ($R^{2}$) and the normalized root mean square error (nRMSE). The nRMSE is computed by normalizing the root mean square error by the range of the traits observations (1–99% quantile), providing a scale-invariant measure of prediction error.
@@ 6 Results and discussions
Table 2: Evaluation of trait prediction with full-range (FR) samples. Trait-wise performance (mean ± standard deviation) includes $R^{2}$ ()($\uparrow$)( ↑ ) and nRMSE ()($\downarrow$)( ↓ ) metrics. Competing methods are: fully supervised baseline (‘Supervised’); MAE with full-range training and linear probing (‘MAE_FR_LP’); MAE with full-range training and fine-tuning (‘MAE_FR_FT’); RTM-based autoencoder (‘RTM_AE’); and semi-supervised regression GAN (‘SR-GAN’). We bold and underline best and second best scores respectively. Trait abbreviations are detailed in Sec. 4.
@@
Table 3: Evaluation of trait prediction with half-range (HR) samples. Trait-wise performance (mean ± standard deviation) includes $R^{2}$ ()($\uparrow$)( ↑ ) and nRMSE ()($\downarrow$)( ↓ ) metrics. Competing methods are: supervised baseline with HR settings (‘Sup_HR’); semi-supervised SR-GAN with HR settings (‘SR-GAN_HR’); RTM-based autoencoder with HR settings (‘RTM-AE_HR’); MAE pretrained on full-range and fine-tuned with linear probing (‘MAE_FR_LP’); MAE pretrained on full-range and fine-tuned (‘MAE_FR_FT’); MAE pretrained on HR and fine-tuned with linear probing (‘MAE_HR_LP’); and MAE pretrained on HR and fine-tuned (‘MAE_HR_FT’). We bold and underline best and second best scores respectively. Trait abbreviations are detailed in Sec. 4.
@@
Trait-specific results, reported with $R^{2}$ and nRMSE scores, are summarized in Tables 2 and 3 for the full- and half-range experiments, respectively. Among all competing methods, the fine-tuned MAE (MAE-FR-FT) outperformed all other methods on most traits when trained and tested on full-range spectra, recording the highest $R^{2}$ values and lowest nRMSE scores. Compared to the fully supervised baseline, MAE-FR-FT led to an average improvement of 9% in $R^{2}$ and 6% in nRMSE. These results underscore the effectiveness of MAEs in learning meaningful spectral representations through masked spectral reconstruction. Pretrained MAE models also exhibited strong cross-spectral generalization, performing competitively on half-range data even when pretrained on full-range spectra and applied to half-range data (MAE-FR-FT). It indicates a good feature transferability and adaptability across heterogeneous sensor configurations, particularly valuable for operational deployment with multi-source data streams.
@@
In the half-range setting, semi- and self-supervised methods also clearly outperformed the supervised baseline. Gains ranged between 100–200% in $R^{2}$ and 17–33% in nRMSE, reinforcing the value of leveraging spectral variability from GreenHyperSpectra even under reduced spectral coverage.
@@
Table 4: Cross-dataset generalization performance. Models are trained on labeled data from all but five datasets (see Sec 4), and evaluated on held-out datasets to assess OOD generalization. Trait-wise performance includes $R^{2}$ ()($\uparrow$)( ↑ ) and nRMSE ()($\downarrow$)( ↓ ) metrics. We bold and underline best and second best scores respectively.
@@
As shown in Table 4, the fine-tuned MAE (MAE-FT) had the highest performance over all other methods across traits, achieving a slight improvement in $R^{2}$ relative to the supervised baseline ($0.31$ vs. $0.24$), along with the lowest average nRMSE. Other competing methods, such as SR-GAN and RTM-AE, provided modest gains over the supervised baseline. Interestingly, MAE with linear probing (MAE-LP) underperformed in the OOD setting, highlighting that fine-tuning is essential to adapt pretrained features to novel spectral distributions. The corresponding scatter plot of the observed and predicted trait values from the different methods is presented in Appendix 16.

### v2

@@ 3 The GreenHyperSpectra dataset
We curated spectral data from multiple instruments across three primary platforms: proximal, airborne, and spaceborne (Figure 2). All data were processed to the level of at-surface reflectance. Proximal measurements were obtained using field spectrometers such as the ASD FieldSpec and SVC HR-1024i, typically positioned in a close range in nadir orientation to record top-of-canopy reflectance. Airborne data were acquired using high-spectral resolution sensors, including the AVIRIS-Next Generation, AVIRIS-Classic, NEON Airborne Observation Platform (AOP) and Specim AISAFenix instruments, which cover landscape-level vegetation scenes with variable viewing geometries and meter-scale spatial resolutions. Spaceborne acquisitions were collected from missions such as PRISMA, Hyperion, EMIT, and EnMAP, offering a larger scale
@@
The dataset includes samples from diverse biomes, with acquisitions spanning from 1992 to 2024, capturing broad ecological and climatic variability across a wide range of environments. Figure 2 maps the global spatial distribution of GreenHyperSpectra and a pool of previously aggregated datasets for plant trait prediction (see details in § 4 ). While the compiled labeled dataset is spatially limited, GreenHyperSpectra encompasses substantially broader spatial coverage and environmental heterogeneity, better representing real-world remote sensing operational conditions (more details in Appendix A).
@@ 5 Experimental settings
We conduct comprehensive ablation experiments across several dimensions to consistently evaluate design trade-offs of the MAE models in spectral representation learning. First, we explore architectural complexity through a grid search spanning transformer configurations with varying numbers of layers $\{6,8,10\}$ and attention heads $\{4,8,16\}$. We select the configuration demonstrating optimal performance on the downstream trait prediction task for subsequent experiments. Considering this optimal architecture, we investigate alternative loss formulations for spectral reconstruction. Beyond conventional MSE, we examine hybrid approaches which incorporate cosine similarity loss weighted by a coefficient $\alpha\in\{1,0.1,0.01,0\}$ to enhance capture of spectral shape specificity. Finally, we assess the effect of token granularity by varying patch sizes (10, 20, 40, and 430) used during spectral masking and reconstruction in the full-range scenario. These targeted ablations informed the final MAE configuration used in our other benchmark evaluations. All results of the aforementioned ablation studies are presented in Tables 16, 17 and 18 in Appendix B.3.
@@
For each experimental setting, we report the performance metrics averaged across three random seeds to measure the variability related to stochastic training effects. Our evaluation framework employs two complementary metrics: the coefficient of determination ($R^{2}$) and the normalized root mean square error (nRMSE). The nRMSE (in %) is computed by normalizing the root mean square error by the range of the traits observations (1–99% quantile), providing a scale-invariant measure of prediction error.
@@ 6 Results and discussions
Table 2: Evaluation of trait prediction with full-range (FR) samples. Trait-wise performance (mean ± standard deviation) includes $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$ metrics. Competing methods are: fully supervised baseline (‘Supervised’); MAE with full-range training and linear probing (‘MAE_FR_LP’); MAE with full-range training and fine-tuning (‘MAE_FR_FT’); RTM-based autoencoder (‘RTM_AE’); and semi-supervised regression GAN (‘SR-GAN’). In RTM-AE, cbc is not directly predicted but is derived from cm and cp estimates (cm – cp). We bold and underline best and second best scores respectively. Trait abbreviations are detailed in Sec. 4.
@@
Table 3: Evaluation of trait prediction with half-range (HR) samples. Trait-wise performance (mean ± standard deviation) includes $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$ metrics. Competing methods are: supervised baseline with HR settings (‘Sup_HR’); semi-supervised SR-GAN with HR settings (‘SR-GAN_HR’); RTM-based autoencoder with HR settings (‘RTM-AE_HR’); MAE pretrained on full-range and fine-tuned with linear probing (‘MAE_FR_HR_LP’); MAE pretrained on full-range and fine-tuned (‘MAE_FR_HR_FT’); MAE pretrained on HR and fine-tuned with linear probing (‘MAE_HR_LP’); and MAE pretrained on HR and fine-tuned (‘MAE_HR_FT’). In RTM-AE, cbc is not directly predicted but is derived from cm and cp estimates (cm – cp). We bold and underline best and second best scores respectively. Trait abbreviations are detailed in Sec. 4.
@@
Trait-specific results, reported with $R^{2}$ and nRMSE scores, are summarized in Tables 2 and 3 for the full- and half-range experiments, respectively. Among all competing methods, the fine-tuned MAE (MAE-FR-FT) outperformed all other methods on most traits when trained and tested on full-range spectra, recording the highest $R^{2}$ values and lowest nRMSE scores. Compared to the fully supervised baseline, MAE-FR-FT led to an average improvement of 9% in $R^{2}$ and 6% in nRMSE. These results underscore the effectiveness of MAEs in learning meaningful spectral representations through masked spectral reconstruction. Pretrained MAE models also exhibited strong cross-spectral generalization, performing competitively on half-range data even when pretrained on full-range spectra and applied to half-range data (MAE-FR-HR-FT). It indicates a good feature transferability and adaptability across heterogeneous sensor configurations, particularly valuable for operational deployment with multi-source data streams.
@@
To further assess robustness of the approaches under sensor noise such as illumination differences, or sensor-specific signal-to-noise characteristics, we additionally evaluated models’ performances under additive Gaussian noise at inference time (details in Tables 26-30). Zero-mean noise with standard deviations of 0.01, 0.03, and 0.05 was added across all spectral bands. Results show that MAE-FR-FT and RTM-AE are substantially more robust than the supervised baseline and GAN. For instance, at $\sigma=0.05$, MAE-FR-FT retains an $R^{2}$ of 0.331 compared to the baseline dropping to –0.065, and both MAE-FR-FT and RTM-AE exhibit smaller increases in nRMSE under spectral corruption, highlighting their resilience.
In the half-range setting, semi- and self-supervised methods also clearly outperformed the supervised baseline. Gains ranged between 100–200% in $R^{2}$ and 8–27% in nRMSE, reinforcing the value of leveraging spectral variability from GreenHyperSpectra even under reduced spectral coverage (Figures  12 and  13).
@@
Table 4: Cross-dataset generalization performance. Models are trained on labeled data from all but five datasets (see Sec 4), and evaluated on held-out datasets to assess OOD generalization. Trait-wise performance includes $R^{2}$ $(\uparrow)$ and nRMSE $(\downarrow)$ metrics. In RTM-AE, cbc is not directly predicted but is derived from cm and cp estimates (cm – cp). We bold and underline best and second best scores respectively.
@@
As shown in Table 4, the fine-tuned MAE (MAE-FR-FT) had the highest performance over all other methods across traits, achieving a slight improvement in $R^{2}$ relative to the supervised baseline ($0.31$ vs. $0.24$), along with the lowest average nRMSE. Since many traits are not associated to single-band features but instead arise from complex interactions across multiple regions of the spectrum, MAE provides a strong prior: it enforces the learning of localized correlations and long-range dependencies within hyperspectral signals, by reconstructing both across adjacent tokens and distant tokens. This prior knowledge facilitates better generalization and more efficient fine-tuning with MAE-FR-FT for the downstream regression. However, this prior alone is not sufficient. When only linear probing is applied (MAE-FR-LP), the model retains general spectral trends leading to underperformance. The necessity of fine-tuning becomes evident in our feature attribution analysis (Fig. 15), where we compared gradient amplitudes across spectral bands for MAE-FR-LP, last block fine-tuning, and full fine-tuning MAE-FR-FT models. While MAE-FR-LP exhibited diffuse and noisy attributions across broad spectral regions, fine-tuning progressively reduced gradient variance, yielding sharper and more interpretable feature importance profiles. This indicates that fine-tuning allows the pretrained prior to be refined toward trait-relevant spectral dependencies, transforming general correlations into targeted representations that drive improved predictive performance.
Other competing methods, such as SR-GAN and RTM-AE, provided modest gains over the supervised baseline. The corresponding scatter plot of the observed and predicted trait values from the different methods is presented in Fig. 14 in the Appendix.
@@ Acknowledgments
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: we summarize four key contributions in Section 1, which are addressed through our experimental design and results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section  7
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This work does not include theoretical analysis.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We share the data and code and provide more details on the experiments in Appendix B.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We submit the code and data URL on the OpenReview platform.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Key training details, including model architecture, loss functions, data splits and major modifications, are summarized in the main text. Comprehensive descriptions of training procedures and hyperparameters are provided in Appendix  B.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We evaluate all models using three random seeds and report the mean and standard deviation of the performance metrics (i.e. $R^{2}$, nRMSE) to reflect variability due to random initialization. This is stated in the main text and reflected in relevant figures and tables (e.g. Figures 5 and  4). For the cross-dataset OOD evaluation, we report results from a single run due to computational constraints; however, we attempt to account for variability through controlled random sampling, as detailed in Section 5.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide detailed information on compute resources in Appendix B, including the type of GPUs, model-specific runtime estimates, and the number of trainable parameters per method (see Table 19).
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS ethical guidelines and ensured that the paper complies with them.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the positive societal impacts of our work in Section  1 and  7. While we did not identify any clear negative impacts, we remain open to addressing potential concerns if raised during the review process.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The data and models presented in this paper do not pose risks of misuse. All datasets are derived from publicly available, non-sensitive sources, and the models are focused on ecological applications.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All datasets used in this work are publicly available and properly credited in the paper. For each source included in GreenHyperSpectra, we provide citation information, access links in Table 7.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The GreenHyperSpectra dataset introduced in this paper is documented in detail in Appendix A. The appendix includes information about data sources, sensor specifications, geographic and temporal coverage, preprocessing steps, and licensing terms for each included dataset. The asset is composed entirely of publicly available data, contains no personal or sensitive information, and respects the original licenses of the sources.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve any crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core methods and experiments presented in this paper do not involve the use of large language models (LLMs).
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Table of figures:
Figure Description
1 Teaser illustration of the proposed semi-/self-supervised frameworks for multi-trait regression.
2 Comparison of GreenHyperSpectra and the labeled dataset, highlighting broader coverage in vegetation types and sensor diversity.
3(a), 3(b) and 3(c) Overview of the SR-GAN, RTM-AE, and MAE architectures for trait prediction.
4 Effect of increasing labeled data volume on R2 performance.
5 Effect of increasing unlabeled pretraining data volume on R2 performance.
6, 7,8 and 9 Dataset characteristics, focusing on spectral variability.
10 and 11 Complementary results to Figs. 4 and  5: nRMSE performance trends with increasing labeled and unlabeled data.
12 and 13 Complementary results to Table 3: Heatmaps of R2 and nRMSE across traits in the half-range input settings to show the performance of MAE vs baseline.
14 Complementary results to Table 4: Observed vs. predicted plots, showing trait-wise calibration in the OOD setting.
15 Feature importance of MAE-based downstream regression as a function of fine-tuning depth (linear probing, final block, and full fine-tuning).
Table 5: Summary of figures and their descriptions.
Table of tables:
Table Description
1 Summary of sensor and platform specifications in GreenHyperSpectra.
2 Trait-wise performance (R2 and nRMSE) of all models under full-range input settings.
3 Trait-wise performance (R2 and nRMSE) of all models under half-range (VNIR) input settings.
4 Trait-wise performance (R2 and nRMSE) of all models under out-of-distribution (OOD) settings.
7 and 8 Dataset details: spectral data characteristics and trait distribution across sources.
9 and 10 Architecture and hyperparameters of SR-GAN.
11, 12 and 13 Architecture, hyperparameters, and RTM configuration of RTM-AE.
14 and 15 Architecture and hyperparameters of MAE.
16, 17 and 18 MAE ablation studies: effects of transformer depth, loss weighting, and token size on trait prediction (R2 and nRMSE).
19 Model size, runtime, and GPU usage across methods.
20, 21, 22, 23, 24 and 25 Complement to Table 4 and Fig. 14: OOD model performance when one vegetation class is excluded from the test set.
26, 27, 28, 29 and 30 Robustness evaluation under additive Gaussian noise during inference, reported for all models (R2 and nRMSE).
Table 6: Summary of tables and their descriptions.